OpenStudy (anonymous):

Michael drove to a friend’s house at a rate of 40 mi/h. He returned by the same route at a rate of 45 mi/h. The driving time for the round trip was 4 h. What is the total distance Michael traveled? a.about 50 miles b.about 169.4 miles c.about 73 miles d.about 201.6 miles

OpenStudy (bahrom7893):

\[d=s*t\]\[d_1=s_1*t_1\]\[d_2=s_2*t_2\]\[d_1=d_2\](since the route was the same), so we now know that: \[s_1*t_1=s_2*t_2\]\[40*t_1=45*t_2\]\[t_1=\frac{45*t_2}{40}=\frac{9*t_2}{8}\]

OpenStudy (bahrom7893):

\[t_1+t_2=4\]\[t_2=4-t_1\]\[t_1=\frac{9*t_2}{8}=\frac{9*(4-t_1)}{8}=\frac{36-9t_1}{8}\]

OpenStudy (bahrom7893):

\[8t_1=36-9t_1\]\[17t_1=36\]\[t_1\approx2.12h\]

OpenStudy (bahrom7893):

\[d_1=s_1*t_1=40*2.12=84.8\]

OpenStudy (bahrom7893):

\[d_2=d_1=84.8\]Total trip is:\[d_1+d_2=84.8+84.8=169.6\]

OpenStudy (bahrom7893):

So, about 169.4 miles

OpenStudy (anonymous):

yep thanks